OpenStudy (anonymous):

A discrete random variable X has a geometric distribution with mean 3. Calculate P(X=5)

OpenStudy (dumbcow):

mean = 1/p -> p = 1/3 P(X=5) = p*(1-p)^(5-1) P(X=5) = (1/3)*(2/3)^4 = 16/243

OpenStudy (anonymous):

Just one quick question.. how did you transfer the mean into a probability?

OpenStudy (dumbcow):

The mean or expected value of a geometric distribution is 1/p

OpenStudy (anonymous):

Ok thanks